a) A beam of 7.8-MeV α particles scatters from a gold foil of thickness 0.30 μm.
i) What fraction of the α particles is scattered between 1.5° and 2.5°?
ii) What is the ratio of α particles scattered through angles greater than 2° to the number
scattered through angles greater than 10°?
1
Expert's answer
2020-06-17T09:18:53-0400

Given, energy of alpha particle "E=7.8\:MeV=12.5\times 10^{-13}J" ,Thickness "t=0.30\mu m=3\times 10^{-7}" .

Let, "Z_1,Z_2" be atomic number of Gold and alpha particle respectively.

"n=\rho_{gold}N_{A}=5.9\times 10^{28} \:atoms/m^3" .


To calculate the number of fractions of alpha particle scattered by angle grater than or equal to "\phi"

is given by,

"F(\phi)=n\pi t\bigg(\frac{Z_1Z_2e^2}{8\pi\epsilon_0E}\bigg)^2\cot^2(\phi/2)"



i).

The fraction of alpha particle scattered between "1.5^{\circ}\leq\phi\leq2.5^{\circ}" is

"F(1.5^{\circ})-F(2.5^{\circ})=0.1176\times10^{-3}\bigg(\cot^2(\frac{1.5^{\circ}}{2})-\cot^2(\frac{2.5^{\circ}}{2})\bigg)\\
F(1.5^{\circ})-F(2.5^{\circ})=0.1176\times10^{-3}\times3735.10\\
F(1.5^{\circ})-F(2.5^{\circ})=0.439"

ii).

Again, from above main formula we get,

"\frac{F(2^{\circ})}{F(10^{\circ})}=\frac{\cot^2(\frac{2^{\circ}}{2})}{\cot^2(\frac{10^{\circ}}{2})}\\
\frac{F(2^{\circ})}{F(10^{\circ})}=\frac{\cot^2(1^{\circ})}{\cot^2(5^{\circ})}\\
\frac{F(2^{\circ})}{F(10^{\circ})}=25.12"